OpenStudy (anonymous):

A right circular cylinder has a base whose area is 154 square feet. If the height of the cylinder is 28 feet, what is its lateral surface area? Use π =

OpenStudy (lgbasallote):

lateral surface area = \(2\pi r h\) does that help?

OpenStudy (anonymous):

Oh alright thanks man! I just didn't have the correct formula.

OpenStudy (lgbasallote):

oh well..you're welcome ^_^

OpenStudy (anonymous):

Oh actually I need one more thing. I need to know the radius. How would I figure that out with only the area?

OpenStudy (lgbasallote):

haha i was typing that out a while ago :p \[A = \pi r^2\] \[\frac{A}{\pi} = r^2\] \[\sqrt{\frac{A}{\pi}} = r\]

OpenStudy (anonymous):

Oh okay haha. Perfect! Thanks a lot for the help!